# Unifying Lane-Level Traffic Prediction from a Graph Structural Perspective: Benchmark and Baseline

### v1

@@ II-C Temporal Dependency Modeling
Time dependency modeling is another crucial aspect of lane-level traffic prediction. The techniques used in existing works for modeling time dependencies can be categorized into machine learning-based methods, RNN-based methods (including those combined with attention mechanisms), and methods based on Graph Neural Networks (GNNs).
@@ II-C1 Machine Learning-based Methods
These methods, such as those using Catboost, Random Forest, and XGBoost [28, 20], rely on traditional statistical and machine learning techniques to analyze time series data. They generally perform well in terms of model interpretability and computational efficiency. However, they may be less effective than deep learning methods in handling complex time dependencies and nonlinear patterns, especially with large-scale or highly dynamic traffic data.
II-C2 RNN-based Methods (Including Attention Mechanisms)
This category includes deep learning models such as LSTM[22, 26, 27], GRU[21, 23, 24, 29], Bi-LSTM[18], and those incorporating attention mechanisms[19]. These methods are effective in capturing long-term dependencies and complex temporal dynamics in traffic flow data. They usually require substantial training time and may need more resources for model training and optimization. Additionally, the complexity of these models can sometimes affect their interpretability.
II-C3 Graph Neural Network-based Methods
Techniques like those utilizing visibility graphs convert the task of learning time dependencies in lane-level traffic flow into mining spatial correlations on a complex spatio-temporal network[25]. This approach offers a novel perspective by transforming time dependencies into visualizable spatial relationships. However, GNNs may not be as directly effective in handling time dependencies as specialized time series models. Moreover, constructing and training GNN models may require more expertise and computational resources.
In summary, while machine learning-based methods perform well in terms of interpretability and computational efficiency, they have limited effectiveness in handling complex temporal dependencies and non-linear patterns. RNN-based methods including models with attention mechanisms, can effectively capture complex temporal dynamics but incur high training costs and their complexity impacts interpretability. Graph Neural Network-based methods offer a novel perspective in processing temporal data, yet they may face challenges in direct effectiveness in modeling temporal dependencies and the resource demands for construction and training.
@@ III Preliminaries and Problem Formulation
Based on the aforementioned literature review, we contend that treating each lane within road segments as an independent modeling object is crucial for enhancing the universality and flexibility of lane-level traffic prediction. As shown in Figure 3, we have appropriately transformed the four spatial topology construction methods to align with lane segments, allowing for flexible usage of individual or combined lane segments to meet diverse spatial topology requirements. In this approach, the lane network is conceptualized as an undirected graph $G=(V,E,A)$, where each vertex $l_{i,j}\in V^{I\times J}$ denotes a lane segment of the $j$th lane of the total $J$ lanes on the $i$th road segments of the total $I$ road segments, where $I\times J=N$, and each edge $e\in E$ indicates whether the lane segments is connected. $A\in\mathbf{R^{N\times N}}$ denotes the static adjacency matrix of graph $G$.
Given a time interval $t$, we use $x_{t}^{l_{i,j}}$ to represent the average traffic state of a lane segment $l_{i,j}$, Towards the whole lane network, we use a state vector, denoted by $X_{t}=[x^{l_{1,1}}_{t},x^{l_{1,2}}_{t},...,x^{l_{I,J}}_{t}]$, to represent the traffic information of all lane segments named supplementary lane segments at the time interval $t$.
@@
were $f$ represents the model used for prediction, and $G$ is an optional, non-essential input.
@@ IV Graph Construction and Simple Baseline
Modeling the lane network as a graph structure provides a structured framework for integrating traffic prediction and modeling at both coarse (such as road-level) and fine (such as lane-level) granularities. This approach not only unifies the spatial topological structures of road-level traffic prediction with lane-level and even city-level predictions, but it also enhances the understanding of complex interactions within the traffic network.
By employing this unified spatial topology, lane-level traffic prediction can be transformed into a spatio-temporal graph prediction problem. This transformation enables the transfer and integration of traffic prediction models from coarse to fine granularities. The graph-based model captures the nuances of lane-specific traffic flows while maintaining a broader perspective of the overall traffic network, allowing for more accurate and comprehensive traffic predictions. Such a unified approach facilitates a deeper analysis of traffic dynamics, providing valuable insights for traffic management and planning.
Figure 4: Overview of GraphMLP model structure and data flow direction.
@@ IV-A Construction of Graph Structure for Lane Network
The adoption of spatio-temporal graph-based algorithms as the mainstream approach in coarse-grained traffic prediction is primarily attributed to their ability to overcome the limitations of Convolutional Neural Networks (CNNs), which are typically confined to Euclidean spaces. spatio-temporal graph models, by contrast, are more adept at processing non-Euclidean structures, such as traffic networks, offering powerful and flexible methods to analyze and predict complex traffic flow patterns. These models provide a fresh perspective for traffic flow prediction by capturing the spatial dependencies and temporal dynamics among network nodes.
In exploring potential coarse-grained models that might be adaptable for fine-grained, lane-level traffic prediction, we consider the following approaches based on spatio-temporal graphs:
@@ IV-A3 Adaptive Graph
IV-B Simple Baseline
When spatial topology is constructed using a graph structure, the possibility of migrating coarse-grained, spatio-temporal graph-based traffic prediction models to lane-level traffic prediction emerges. Despite showing potential, these models in their original design did not fully consider the unique challenges faced by lane-level prediction. The complexity of lane-level traffic prediction is primarily manifested in more detailed and complex spatial dependencies, high demands for real-time data processing, and the need for dynamic predictive capabilities.
@@ IV-B2 Instance Normalization
$\hat{x}^{l_{i,j}}=\psi\big{(}\frac{x^{l_{i,j}}-\mu(x^{l_{i,j}})}{\sqrt{\sigma^{2}(x^{l_{i,j}})}}\big{)}+\beta$ (5)
@@ IV-B3 Dynamic Graph Network
Due to the spatial complexity of lane-level road networks, the configuration of adjacent lane segments may present in multiple directions, such as up, down, left, right, and their diagonal counterparts. This complex spatial arrangement can make traditional fixed-contact adjacency matrices less flexible in capturing relationships between lanes. Therefore, to more accurately reflect the spatial relationships between lane segments, we adopted a self-attention mechanism to dynamically identify adjacent lanes relevant to each lane segment. This approach allows the adjacency matrix to change dynamically over time, better capturing the complex and time-varying spatial connections between lanes. The self-attention mechanism generates an adjacency matrix by calculating the relationships between each node and all other nodes. Given the traffic state vector $x^{l_{i,j}}$ of lane segment $l_{i,j}$, we first compute the Query ($Q$), Key ($K$), and Value ($V$) using learnable weight matrices:
$Q^{l_{i,j}}=W^{Q}x^{l_{i,j}}$ (6)
$K^{l_{i,j}}=W^{K}x^{l_{i,j}}$
$V^{l_{i,j}}=W^{V}x^{l_{i,j}}$
Where $W^{Q},W^{K},W^{V}$ are the weight matrices for Query, Key, and Value, respectively. The attention coefficient $\alpha_{i_{i,j},l_{a,b}}$ between lane segment $l_{i,j}$ and lane segment $l_{a,b}$ is then computed as:
@@
In this equation, $d_{k}$ is the dimension of the key vector, which scales the dot product to avoid overly large values. The softmax function ensures that the sum of all output attention coefficients for a given lane segment $l_{i,j}$ equals 1.
Using this dynamic adjacency matrix, we further perform graph convolution operations, effectively combining the spatial information of adjacent lane segments. the new feature of each node is computed by aggregating the weighted sum of its own features and the features of its neighboring nodes. The new feature $\bar{x}^{l_{i,j}}$ of lane segment $l_{i,j}$ is calculated as follows:
$\bar{x}^{l_{i,j}}=\rho\big{(}\sum_{l_{a,b}\in\mathcal{N}(l_{i,j})}\alpha_{l_{i,j},l_{a,b}}x^{l_{a,b}}\big{)}$ (8)
@@ IV-B4 Independent Temporal MLP Network
Although various Transformer model variants [34, 35, 36] have been adapted to suit the characteristics of time series data, it has been demonstrated that a remarkably simple linear model, DLinear[37], can outperform most of these Transformer-based forecasting models. The current research provides a solid foundation for designing a simple yet efficient Independent MLP (Multilayer Perceptron) network layer suitable for temporal-dependency modeling in lane-level traffic networks.
Initially, we segment the state vector $x^{l_{i,j}}$ of each individual lane segment into equally-sized patches $p$:
$p_{n}=x^{l_{i,j}}_{(n-1)k:nk},\text{for }n=1,...,m$ (9)
@@ IV-B6 Instance Denormalizatiom
Here, $\theta$ represents the ensemble of adjustable parameters within the GraphMLP framework. This approach ensures a rigorous adaptation of the model to the data, striving for minimal deviation between predicted and actual values.
@@ V-A Datasets
We have collected and processed two real-world lane-level traffic datasets, including data from the well-known public Performance Measurement System (PeMS [^1] https://pems.dot.ca.gov/ ) for Los Angeles’ Santa Ana Freeway and urban expressway data from sensors on the HuaNan City Expressway in Guangzhou, China. Based on these real datasets, we generated three datasets: the PeMS dataset (with 5 lanes) and the HuaNan dataset (with 4 lanes) representing regular lane datasets, while the PeMSF dataset includes entrance lanes, meaning some road segments contain a sixth lane, representing an irregular lane count dataset. Figure 5 displays the sensor locations in both data sources and the segments containing entrance lanes.
Evaluating the model on these two types of real datasets, we aim to assess its effectiveness, accuracy, and versatility in predicting lane-level traffic under various conditions and complexities. This comprehensive approach ensures that the model can be tested not only in controlled or limited scenarios but also in a broader range of real-world traffic situations.

### v2

@@ II-C Temporal Dependency Modeling
Temporal dependency modeling is another crucial component of lane-level traffic prediction, aiming to capture the dynamic evolution of traffic flow over time. This process involves characterizing not only short-term fluctuations but also long-term temporal dependencies and potential nonlinear variations. Existing methods for temporal modeling differ significantly in their technical approaches and can be broadly categorized into machine learning-based methods, RNN-based deep learning methods (including attention-enhanced variants), and emerging GNN-based methods. Below, these approaches are summarized and evaluated from the perspectives of theoretical implementation and practical effectiveness.
@@ II-C1 Machine Learning-based Methods
Machine learning-based methods rely on traditional statistical and machine learning techniques to predict traffic flow by extracting temporal features (e.g., time-of-day effects, trends, and periodicity) through feature engineering. Common approaches include CatBoost, Random Forest (RF), and Gradient Boosting Trees (XGBoost)[20, 21]. Compared to deep learning models, these methods excel in computational efficiency and model interpretability, making them suitable for scenarios with small datasets or relatively simple temporal dependencies. For instance, analyzing the importance of features enables a clear explanation of the time-driven factors that influence traffic flow.
However, these methods are constrained by their linear assumptions and limited feature representation capabilities. They struggle to handle complex temporal dependencies and nonlinear dynamic variations, particularly in lane-level traffic prediction. These limitations make them inadequate for capturing long-term correlations in traffic flow or responding effectively to high-dynamic changes caused by sudden events, such as traffic accidents.
II-C2 RNN-based Methods
RNNs and their extensions are core technologies in time series modeling and have been widely applied in lane-level traffic prediction in recent years. Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRU) effectively capture both short-term and long-term dependencies in time series through the use of gating mechanisms[23, 24, 25, 28, 26, 27, 29]. For example, Bidirectional LSTM (Bi-LSTM) enhances the modeling of traffic flow dynamics by considering both historical and future time segments via bidirectional information flow[19].
Attention-enhanced temporal modeling methods further improve model flexibility by dynamically assigning weights to focus on the most critical time segments for prediction. These approaches are particularly effective in capturing traffic surges caused by short-term events, such as traffic signal changes. However, these methods face several challenges. First, the sequential training mechanism of RNNs and their extensions results in lower computational efficiency, especially for long-time series, making the training process time-intensive. Second, the increased model complexity raises resource demands and limits scalability for large-scale lane flow data. While attention mechanisms improve interpretability to some extent, the hidden states within RNNs make the overall model difficult to interpret comprehensively [34, 35].
II-C3 GNN-based Methods
GNN-based methods offer an innovative approach to temporal dependency modeling. Unlike traditional time series methods, these approaches transform time series into graph structures, representing time points as graph nodes and temporal dependencies as edges between nodes. For example, the Visibility Graph technique constructs a visualized spatial relation network from time series data and leverages GCNs to learn dependency relationships[22].
Compared to traditional methods, GNN-based approaches can unify spatial and temporal information, which is particularly beneficial for lane-level scenarios where both dimensions interact closely. However, the efficiency of GNN-based temporal modeling remains a concern. Transforming temporal dependencies into graph relationships may lead to suboptimal performance in some complex time series tasks compared to specialized time series models. Additionally, constructing dynamic temporal graphs requires advanced modeling techniques and significant computational resources, raising the application threshold for these methods.
II-C4 Comprehensive Analysis
Temporal dependency modeling methods have distinct characteristics, and their selection typically depends on the specific prediction task and data properties. Machine learning-based methods excel in computational efficiency and interpretability but are limited in their adaptability to complex temporal dynamics. RNN-based methods, particularly those enhanced with attention mechanisms, leverage deep networks to capture intricate temporal dependencies, making them suitable for dynamic traffic data with long-term correlations. In contrast, GNN-based methods introduce a spatio-temporal unified modeling perspective through graph structures, offering a novel approach to handling complex interactions in lane-level traffic flows.
In practical applications, selecting a method requires balancing prediction accuracy, computational efficiency, and modeling complexity. Future research should focus on exploring more efficient deep learning architectures, particularly in lightweight modeling and multimodal data integration, to develop more accurate and efficient temporal dependency modeling solutions for lane-level traffic prediction.
@@ III Preliminaries and Problem Formulation
Based on the aforementioned literature review, we contend that treating each lane within road segments as an independent modeling object is crucial for enhancing the universality and flexibility of lane-level traffic prediction. As shown in Figure 3, we have appropriately transformed the four spatial topology construction methods to align with lane segments, allowing for flexible usage of individual or combined lane segments to meet diverse spatial topology requirements. In this approach, the lane network is conceptualized as an undirected graph $G=(V,E,A)$, where each vertex $l_{i,j}\in V^{N}$ denotes a lane segment of the $j$th lane of the total $J_{i}$ lanes on the $i$th road segments of the total $I$ road segments, where $\sum_{i=0}^{I}J_{i}=N$, and each edge $e\in E$ indicates whether the lane segments is connected.$A\in\mathbf{R^{N\times N}}$ denotes the static adjacency matrix of graph $G$. Furthermore, adjacency relationships are defined by both sequential lane connections and permitted lane changes, meaning adjacent lanes are not connected where lane changes are prohibited. At intersections, connectivity is determined by valid turning movements, ensuring an accurate representation of spatial dependencies. This graph-based formulation not only facilitates accurate prediction but also enables mining interpretable patterns and dependencies across lane segments.
Given a time interval $t$, we use $x_{t}^{l_{i,j}}$ to represent the average traffic state of a lane segment $l_{i,j}$, Towards the whole lane network, we use a state vector, denoted by $X_{t}=[x^{l_{1,1}}_{t},x^{l_{1,2}}_{t},...,x^{l_{I,J_{I}}}_{t}]$, to represent the traffic information of all lane segments named supplementary lane segments at the time interval $t$.
@@
where $f$ represents the model used for prediction, and $G$ is an optional, non-essential input.
@@ IV Graph Construction and Simple Baseline
We propose a structured framework to model various lane networks as a unified graph structure, integrating coarse-grained (e.g., road-level) and fine-grained (e.g., lane-level) traffic prediction and modeling. This approach not only unifies the spatial topology of road-level traffic prediction with lane-level and even city-level predictions but also enhances the understanding of complex interactions within traffic networks. Utilizing this unified spatial topology, lane-level traffic prediction can be formulated as a spatiotemporal graph prediction problem. This transformation enables seamless integration and conversion of traffic prediction models across different granularities, from coarse to fine.
Graph-based models capture the nuances of traffic flow on specific lanes while maintaining a broader perspective of the overall traffic network, thereby achieving more accurate and comprehensive traffic predictions. This unified approach facilitates deeper analysis of traffic dynamics, offering valuable insights for traffic management and planning.
Figure 4: Overview of GraphMLP model structure and data flow direction. $\mu$ and $\sigma$ represent the mean and standard deviation of lane segments, respectively.
@@ IV-A Construction of Graph Structure for Lane Network
The adoption of spatio-temporal graph-based algorithms as the mainstream approach in coarse-grained traffic prediction is primarily attributed to their ability to overcome the limitations of CNNs, which are typically confined to Euclidean spaces. spatio-temporal graph models, by contrast, are more adept at processing non-Euclidean structures, such as traffic networks, offering powerful and flexible methods to analyze and predict complex traffic flow patterns. These models provide a fresh perspective for traffic flow prediction by capturing the spatial dependencies and temporal dynamics among network nodes.
As illustrated in Figure 3, we demonstrate the segmentation of lane networks to encompass all existing topological structures studied. In exploring potential coarse-grained models that might be adaptable for fine-grained, lane-level traffic prediction, we consider the following approaches based on spatio-temporal graphs:
@@ IV-A3 Adaptive Graph
IV-B A Simple but Effective New Baseline
The aforementioned graph-based modeling approach not only offers a graph-structured alternative design for existing models but also can be applied to resource-constrained edge devices. Despite its potential, it does not fully address the unique challenges of lane-level prediction. The complexity of lane-level traffic prediction is primarily manifested in more detailed and complex spatial dependencies, high demands for real-time data processing, and the need for dynamic predictive capabilities.
@@ IV-B2 Instance Normalization
$\dot{x}^{l_{i,j}}=\psi\big{(}\frac{x^{l_{i,j}}-\mu(x^{l_{i,j}})}{\sqrt{\sigma^{2}(x^{l_{i,j}})}}\big{)}+\beta$ (5)
@@ IV-B3 Dynamic Graph Network
Due to the spatial complexity of lane-level road networks, the configuration of adjacent lane segments may present in multiple directions, such as up, down, left, right, and their diagonal counterparts. This complex spatial arrangement can make traditional fixed-contact adjacency matrices less flexible in capturing relationships between lanes. Therefore, to more accurately reflect the spatial relationships between lane segments, we adopted a self-attention mechanism to dynamically identify adjacent lanes relevant to each lane segment. This approach allows the adjacency matrix to change dynamically over time, better capturing the complex and time-varying spatial connections between lanes. The self-attention mechanism generates an adjacency matrix by calculating the relationships between each node and all other nodes. Given the traffic state vector $\dot{x}^{l_{i,j}}$ of lane segment $l_{i,j}$, we first compute the Query ($Q$), Key ($K$), and Value ($V$) using learnable weight matrices:
$Q^{l_{i,j}}=W^{Q}\dot{x}^{l_{i,j}}$ (6)
$K^{l_{i,j}}=W^{K}\dot{x}^{l_{i,j}}$
$V^{l_{i,j}}=W^{V}\dot{x}^{l_{i,j}}$
where $W^{Q},W^{K},W^{V}$ are the weight matrices for Query, Key, and Value, respectively. These weight matrices are trainable parameters that allow the model to learn specific patterns and relationships between different traffic states during training. The attention coefficient $\alpha_{i_{i,j},l_{a,b}}$ between lane segment $l_{i,j}$ and lane segment $l_{a,b}$ is then computed as:
@@
In this equation, $d_{k}$ is the dimension of the key vector, which scales the dot product to avoid overly large values. The softmax function ensures that the sum of all output attention coefficients for a given lane segment $l_{i,j}$ equals 1. This normalization ensures that the relative importance of each neighboring lane segment is appropriately weighted, maintaining the consistency and interpretability of the attention mechanism.
Using this dynamic adjacency matrix, we further perform graph convolution operations, effectively combining the spatial information of adjacent lane segments. The integration of self-attention coefficients into the adjacency matrix allows the model to dynamically adapt to traffic patterns by prioritizing critical spatial dependencies. The new feature of each node is computed by aggregating the weighted sum of its own features and the features of its neighboring nodes. The new feature $\bar{x}^{l_{i,j}}$ of lane segment $l_{i,j}$ is calculated as follows:
$\bar{x}^{l_{i,j}}=\rho\big{(}\sum_{l_{a,b}\in\mathcal{N}(l_{i,j})}\alpha_{l_{i,j},l_{a,b}}\dot{x}^{l_{a,b}}\big{)}$ (8)
@@ IV-B4 Independent Temporal MLP Network
Although various Transformer model variants [40, 41, 42] have been adapted to suit the characteristics of time series data, it has been demonstrated that a remarkably simple linear model, DLinear[43], can outperform most of these Transformer-based prediction models. The current research provides a solid foundation for designing a simple yet efficient Independent MLP (Multilayer Perceptron) network layer suitable for temporal-dependency modeling in lane-level traffic networks.
Initially, we segment the state vector $\dot{x}^{l_{i,j}}$ of each individual lane segment into equally-sized patches $p$:
$p_{n}=\dot{x}^{l_{i,j}}_{(n-1)k:nk},\text{for }n=1,...,m$ (9)
@@ IV-B5 Gating Network
Input: Graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, feature matrix $\mathbf{X}$, ground truth labels $\mathbf{Y}$, number of epochs $T_{train}$, learning rate $\eta$, loss function $\mathcal{L}$.
Output: Trained parameters $\theta$.
Step 1: Model Initialization. Initialize model parameters $\theta$, including the dynamic graph module, independent temporal MLP, gating network, and normalization parameters.
for epoch $t=1$ to $T_{train}$ do
      Step 2: Instance Normalization. Normalize the input feature $\mathbf{X}$ for each lane segment $l_{i,j}$ using:
$\dot{x}_{l_{i,j}}=\psi\cdot\frac{x_{l_{i,j}}-\mu(x_{l_{i,j}})}{\sqrt{\sigma^{2}(x_{l_{i,j}})}}+\beta$
where $\mu(x_{l_{i,j}})$ and $\sigma^{2}(x_{l_{i,j}})$ are the mean and variance over the time window for $l_{i,j}$, and $\psi$, $\beta$ are learnable parameters.
      Step 3: Spatial Dependency Modeling. Construct the dynamic adjacency matrix and compute attention coefficients:
$\alpha_{l_{i,j},l_{a,b}}=\text{Softmax}\Big{(}\text{LeakyReLU}\Big{(}\frac{Q_{l_{i,j}}\cdot K_{l_{a,b}}^{\top}}{\sqrt{d_{k}}}\Big{)}\Big{)}$
Perform graph convolution to update spatial features:
$\bar{\mathbf{X}}=\rho\Big{(}\sum_{l_{a,b}\in\mathcal{N}(l_{i,j})}\alpha_{l_{i,j},l_{a,b}}\cdot\dot{\mathbf{X}}\Big{)}$
      Step 4: Temporal Dependency Modeling. Divide the normalized temporal input into patches:
$p_{n}=\dot{\mathbf{X}}[(n-1)k:nk],\quad n=1,\dots,m$
Process patches through the independent temporal MLP:
$p^{\prime}_{n}=\text{MLP}(p_{n}),\quad\hat{\mathbf{X}}=\text{MLP}(p^{\prime}_{1},\dots,p^{\prime}_{m})$
      Step 5: Gating Network for Feature Fusion. Fuse spatial and temporal features:
$\bar{\mathbf{Y}}=\lambda\cdot\bar{\mathbf{X}}+(1-\lambda)\cdot\hat{\mathbf{X}}$
where $\lambda$ is the gating weight.
      Step 6: Instance Denormalization. Restore the statistical properties removed during normalization for each lane segment $l_{i,j}$ in $\hat{Y}$ using:
$\hat{y}^{l_{i,j}}=\sqrt{\sigma(x^{l_{i,j}})}\left(\frac{\bar{y}^{l_{i,j}}-\beta}{\psi}\right)+\mu(x^{l_{i,j}})$
      Step 7: Loss Computation and Backpropagation. Compute loss:
$\mathcal{L}=\text{Loss}(\hat{\mathbf{Y}},\mathbf{Y})$
Update parameters:
$\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}$
end for
return $\theta$;
Algorithm 1 Training Procedure for GraphMLP
@@ IV-B6 Instance Denormalizatiom
where $\theta$ represents the ensemble of trainable parameters within the GraphMLP framework. Algorithm 1 presents the pseudocode for the training process of GraphMLP. GraphMLP offers a more lightweight and adaptable foundational solution for lane-level traffic prediction, capable of accommodating the spatial complexity and dynamics of lanes.
@@ V Benchmark
In this benchmark, we first introduce three benchmark datasets that we have collected, processed, and made publicly available. These datasets represent regular and irregular freeways, namely the PeMS and PeMSF datasets, as well as the HuaNan dataset representing urban expressways. By treating lanes as a unified graph structure and utilizing the adjacency relationships constructed as described in Section IV, we introduce spatio-temporal graph models to advance lane-level traffic prediction. Additionally, we conduct a unified experimental comparison of these models with lane-level prediction algorithms, providing more challenging benchmarks for the field of lane-level traffic prediction.
@@ V-A Datasets
We have collected and processed real-world lane-level traffic data from two representative sources: the public Performance Measurement System (PeMS [^1] https://pems.dot.ca.gov/) for the Santa Ana Freeway in Los Angeles, and sensor data from the HuaNan Expressway in Guangzhou, China. Based on these two sources, we constructed three distinct datasets to support our benchmark experiments. The PeMS dataset features a regular 5-lane configuration. The HuaNan dataset captures a typical 4-lane expressway in an urban environment. In contrast, the PeMSF dataset is an extended variant of PeMS that includes freeway segments with entrance ramps, where some road segments contain a sixth lane, representing irregular lane configurations.
This triplet of datasets is deliberately designed to encompass both regular and irregular lane conditions, enabling comprehensive evaluation of lane-level traffic prediction models. Figure 5 displays the sensor locations across both sources, along with examples of segments containing entrance lanes.
Evaluating the model on these datasets allows us to assess its effectiveness, accuracy, and generalizability across varying lane structures and real-world complexities. This ensures that the benchmark is not confined to simplified or idealized road scenarios, but rather reflects a wide spectrum of practical traffic conditions.